OpenStudy (anonymous):

Help............275^4=36^7/9x^21

OpenStudy (anonymous):

I have to post an answer in 3 minutes!!!!!!!

OpenStudy (anonymous):

\[9x^{21}=\frac{ 36^7 }{ 275^4 }\] \[x^{21}=\frac{ 36^7 }{ 9(275)^4 }\]